On December 31, 2015, Culver Co. performed environmental consulting services for Hayduke Co. Hayduke was short of cash, and Culver Co. agreed to accept a $318,100 zero-interest-bearing note due December 31, 2017, as payment in full. Hayduke is somewhat of a credit risk and typically borrows funds at a rate of 11%. Culver is much more creditworthy and has various lines of credit at 5%.

Prepare the journal entry to record the transaction of December 31, 2015, for the Culver Co.
Answer #1

First we will have to calculate the service revenue earned by Culver.

Service revenue earned by Culver in 2015 = amount of notes receivable*PVIF (11%, 2)

Using the present value interest factor (PVIF) table we find that for 11% and 2 years the applicable discount factor = 0.8116

Thus revenue = 318,100*0.8116 = 258,170

Difference between the note value and service revenue will be the discount amount. Hence discount = 318,100 - 258,170 = 59,930

Thus journal entry for Culver:

Date Particulars Debit Credit
12/31/2015 Notes receivable 318,100
Discount on notes receivable 59,930
Service revenue 258,170